# Note on Regularity theory for Nonlinear elliptic equations

### v2

@@ 2.2. DNM theory I: De Giorgi-Moser iteration
$\left(\int_{B_{1}}(\varphi_{k}v_{k})^{p}\right)^{2/p}\lesssim\int_{B_{1}}|\nabla(\varphi_{k}v_{k})|^{2}.$
@@
$\int_{B_{1}}(\varphi_{k}v_{k})^{2}\lesssim\left(\int_{B_{1}}(\varphi_{k}v_{k})^{p}\right)^{2/p}\cdot|\{\varphi_{k}v_{k}>0\}|^{\epsilon},$
@@
$A_{k}\lesssim|\{\varphi_{k}v_{k}>0\}|^{\epsilon}\cdot\int_{B_{1}}|\nabla(\varphi_{k}v_{k})|^{2}=:|\{\varphi_{k}v_{k}>0\}|^{\epsilon}\cdot E_{k}.$
@@ 1. Step 1
    (The case $p\geq 2$) For a non‑negative test function $\varphi\in H^{1}_{0}(B_{1})$ we have
@@
    $\int_{B_{1}}|\nabla v|^{2}v^{p-2}\eta^{2}\leq C\int_{B_{1}}|\nabla v||\nabla\eta|\,\eta v^{p-1}.$
    Rewrite the right‑hand side as
    $\int_{B_{1}}\left(|\nabla v|v^{\frac{p-2}{2}}\eta\right)\left(|\nabla\eta|v^{\frac{p}{2}}\right)$
    and apply Cauchy inequality $ab\leq\frac{\varepsilon}{2}a^{2}+\frac{1}{2\varepsilon}b^{2}$ for suitable $\epsilon$, we get
    $\int_{B_{1}}|\nabla v|^{2}v^{p-2}\eta^{2}\leq C\int_{B_{1}}|\nabla\eta|^{2}v^{p}.$
@@
    (2.9) $\int_{B_{1}}|\nabla(v^{p/2})|^{2}\eta^{2}\leq C\int_{B_{1}}|\nabla\eta|^{2}v^{p},$
    where $C$ is up to $p,\lambda,\Lambda$.
@@ 2. Step 2
    $\left(\int_{B_{1}}(\eta v^{p/2})^{\frac{2n}{n-2}}\right)^{\!\frac{n-2}{n}}\leq C_{S}\int_{B_{1}}\left(|\nabla(\eta v^{p/2})|^{2}+|\nabla\eta|^{2}v^{p}\right).$
    Using $|\nabla(\eta v^{p/2})|\leq|\nabla\eta|v^{p/2}+\eta|\nabla(v^{p/2})|$ together with (2.9) we obtain
    (2.10) $\left(\int_{B_{1}}\eta^{\frac{2n}{n-2}}v^{\frac{np}{n-2}}\right)^{\!\frac{n-2}{n}}\leq\tilde{C}\int_{B_{1}}|\nabla\eta|^{2}v^{p},$
    where $\tilde{C}=CC_{S}$ depends on $p,n,\lambda,\Lambda$.
@@ 3. Step 3
    $\eta_{k}:=\begin{cases}1,&\text{in}\ B_{1/2+2^{-k}},\\ 0,&\text{outside}\ B_{1/2+2^{-k+1}}.\end{cases}$
@@
    $\left(\int_{B_{r_{1/2+2^{-k-1}}}}v^{p_{k+1}}\right)^{\!1/p_{k+1}}\leq\left(\tilde{C}2^{2k}\right)^{1/p_{k}}\left(\int_{B_{1/2+2^{-k}}}v^{p_{k}}\right)^{\!1/p_{k}}.$
    Iterating this estimate from $k=0$ to $k=m$ gives
    $\|v\|_{L^{p_{m+1}}(B_{1/2+{2^{-m-1}}})}\leq\left(\prod_{k=0}^{m}(\tilde{C}2^{2k})^{1/p_{k}}\right)\|v\|_{L^{p}(B_{1})}.$
    Since $\sum_{k}1/p_{k}<\infty$ and $\sum_{k}2k/p_{k}<\infty$, the product converges to a finite constant $C_{0}=C_{0}(n,\lambda,\Lambda,p)$. Letting $m\to\infty$ we obtain
@@ 4. Step 4
    (The case $0<p<2$) For any $\delta>0$, write
    $v=v\chi_{\{v\leq\delta\}}+v\chi_{\{v>\delta\}}\leq\delta+v\chi_{\{v>\delta\}}.$
    Then
    $\|v\|_{L^{2}(B_{1})}^{2}\leq 2\delta^{2}|B_{1}|+2\int_{B_{1}}v^{2}\chi_{\{v>\delta\}}\leq 2\delta^{2}|B_{1}|+2\int_{B_{1}}v^{2}\left(\frac{v}{\delta}\right)^{2-p}=2\delta^{2}|B_{1}|+2\delta^{p-2}\int_{B_{1}}v^{p}.$
    Choosing $\delta=\left(\int_{B_{1}}v^{p}\right)^{1/p}$, we obtain
    (2.12) $\|v\|_{L^{2}(B_{1})}\lesssim\|v\|_{L^{p}(B_{1})}.$
    Combining (2.11) and (2.12) yields
    $\|v\|_{L^{\infty}(B_{1/2})}\lesssim\|v\|_{L^{p}(B_{1})},$
    which completes the proof for $0<p<2$.
@@
(2.13) $\inf_{B_{1/2}}v\gtrsim\left(\int_{B_{1}}v^{-p}\right)^{\!-1/p}.$
@@
$\int_{B_{1}}e^{p_{0}|w-\bar{w}_{B_{1}}|}\leq C,$
where $\bar{w}_{B_{1}}=\frac{1}{|B_{1}|}\int_{B_{1}}w$. Take $p=p_{0}$ in (2.13). ∎
@@ 3.1. Introduction to fully nonlinear equations
Observe that the $L^{\epsilon}$-norm of a superharmonic function is controlled by its value at a point; hence, if we fix the value of $v$ at some point, say $v(0)=1$ without loss of generality, then $v$ belongs to $L^{\epsilon}$. See Lemma 4
Conversely, the value of a subharmonic function is controlled by the nearby $L^{\epsilon}$-norm, so a fixed $L^{\epsilon}$-norm may govern the value at the center. See Lemma 1 and Lemma 3.
@@ 3.3. Krylov-Safonov weak Harnack inequality
Let $v$ be a supersolution of (3.4) with $v(0)=1$. Then $v\in L^{\epsilon}_{weak}(B_{1/2})$, i.e.,
$|\{v>\lambda\}|\leq\lambda^{-\epsilon},\quad\forall\lambda>0.$
@@
First, we derive a measure estimate via the ABP method. We claim that
@@
If $x_{0}\in\Gamma^{+}$, then $\tilde{v}(x_{0})<0$ (noting that $\tilde{v}|_{\partial B_{1}}\geq 0$), which implies $v(x_{0})<2$. Hence $\Gamma^{+}\subset\{v<2\}$, and therefore
@@
sublemma 4 (Measure estimate).
@@
with constants chosen so that
@@
and set $\tilde{v}:=v+\gamma$ (see Figure 9). We can arrange $\gamma$ so that $\tilde{v}(0)=-2$.
@@
If $x\in\Gamma^{+}$, then $\tilde{v}(x)<0$, which implies $v(x)<M$ for a suitable choice of $M$. Hence $\Gamma^{+}\subset\{v<M\}\cap B_{1/8}(x_{0})$, and the desired inequality follows. ∎
@@
which also solves (3.4). We consider two cases; if the first does not hold, we replace $w$ by 1w1-w.
@@ 4.1. Concave assumption of $F$ and Gap
It remains to prove that $u\in C^{2,\alpha}$, i.e., $D^{2}u\in C^{\alpha}$. The $C^{\alpha}$ estimate (Corollary 4) requires a solution, but we only know that $v$ is a subsolution. This gap constitutes the main difficulty in the Evans-Krylov theory. However, the relation
@@ 4.2. Deletion argument
Cover $\mathcal{H}(B_{1})$ by balls of radius $\delta$; their number satisfies $\lesssim\delta^{-n^{2}}$. Hence some ball $B_{\delta}(M_{3})$ satisfies the measure estimate
@@
For any $z\in\mathcal{H}^{-1}(B_{\delta}(M_{3}))$, we have $|u_{gg}(z)-u_{gg}(x_{3})|\leq C\delta$. Choosing $\delta$ sufficiently small (universally) gives
$\mathbf{t}-u_{gg}(z)\gtrsim\frac{1}{8}-C\delta\gtrsim\frac{1}{8},$
@@ 4.3.1. Krylov-Safonov
A simpler proof of the Krylov-Safonov theorem can be found in [mooney2019proof]. The ideas behind their Harnack inequality were further developed by Caffarelli [caffarelli1987harnack] and Caffarelli-Wang [caffarelli1993harnack], and later extended by Savin [savin2007small] to nearly flat solutions (small perturbation).
4.3.2. Evans-Krylov
Since the Evans-Krylov theory assumes concavity of $F$, a natural question is whether solutions are always $C^{2}$. The answer is affirmative for $n=2$ [nirenberg1953nonlinear] (see also Theorem 4.8 of [fernandez2023regularity]), where solutions are even $C^{2,\alpha}$. Counterexamples for $n\geq 5$ were constructed by Nadirashvili-Vladuts [nadirashvili2007nonclassical, nadirashvili2008singular, nadirashvili2013singular]. The cases $n=3,4$ remain open. We refer to [caffarelli1995fully] for a comprehensive treatment of fully nonlinear equations.

### v3

@@ 2.2. DNM theory I: De Giorgi-Moser iteration
$\left(\int(\varphi_{k}v_{k})^{p}\right)^{2/p}\lesssim\int|\nabla(\varphi_{k}v_{k})|^{2}.$
@@
$\int(\varphi_{k}v_{k})^{2}\lesssim\left(\int(\varphi_{k}v_{k})^{p}\right)^{2/p}\cdot|\{\varphi_{k}v_{k}>0\}|^{\epsilon},$
@@
$A_{k}\lesssim|\{\varphi_{k}v_{k}>0\}|^{\epsilon}\cdot\int|\nabla(\varphi_{k}v_{k})|^{2}=:|\{\varphi_{k}v_{k}>0\}|^{\epsilon}\cdot E_{k}.$
@@ 1. Step 1
    (The case $p\geq 2$) For a nonnegative test function $\varphi\in H^{1}_{0}(B_{1})$ we have
@@
    $\int|\nabla v|^{2}v^{p-2}\eta^{2}\leq C\int|\nabla v||\nabla\eta|\,\eta v^{p-1}.$
    Rewrite the right‑hand side as $\int\left(|\nabla v|v^{\frac{p-2}{2}}\eta\right)\left(|\nabla\eta|v^{\frac{p}{2}}\right)$, and apply Young inequality $ab\leq\frac{\varepsilon}{2}a^{2}+\frac{1}{2\varepsilon}b^{2}$, we obtain
    $\int|\nabla v|^{2}v^{p-2}\eta^{2}\leq C_{\epsilon}\int|\nabla\eta|^{2}v^{p}.$
@@
    (2.9) $\int|\nabla(v^{p/2})|^{2}\eta^{2}\leq C_{1}\int|\nabla\eta|^{2}v^{p},$
    where $C_{1}$ is up to $\epsilon,p,\lambda,\Lambda$.
@@ 2. Step 2
    $\left(\int(\eta v^{p/2})^{\frac{2n}{n-2}}\right)^{\!\frac{n-2}{n}}\leq C_{S}\int\left(|\nabla\eta|v^{p/2}+\eta|\nabla(v^{p/2})|\right)^{2}.$
    Combining with (2.9) and using Young inequality again, we obtain
    (2.10) $\left(\int v^{\frac{n}{n-2}\cdot p}\cdot\eta^{\frac{2n}{n-2}}\right)^{\frac{n-2}{n}}\leq C_{2}\int|\nabla\eta|^{2}v^{p},$
    where $C_{2}$ depends on $p,n,\lambda,\Lambda$ for sufficiently small $\epsilon$.
@@ 3. Step 3
    $\eta_{k}:=\begin{cases}1,&\text{in}\ B_{1/2+2^{-k}},\\ 0,&\text{outside}\ B_{1/2+2^{-k+1}}\end{cases},\quad k=1,2,\cdots.$
@@
    $\int_{B_{1/2+2^{-k}}}v^{p_{k}}\lesssim C2^{2k(1+\epsilon)}\left(\int_{B_{1/2+2^{-k+1}}}v^{p_{k-1}}\right)^{1+\epsilon},\quad\epsilon=\frac{2}{n-2}>0,$
    which leads to the same argument in the proof of Lemma 1. Thus, we obtain
@@ 4. Step 4
    (The case $0<p<2$) By (2.11) with $p=2$ and Young inequality, we obtain
    $\sup_{B_{1/2}}v\lesssim\epsilon\sup_{B_{1}}v^{1-\frac{p}{2}}+C_{\epsilon}\left(\int v^{p}\right)^{1/2}.$
    Finally, by absorbing lemma, we complete the proof.
@@
(2.12) $\inf_{B_{1/2}}v\gtrsim\left(\int v^{-p}\right)^{\!-1/p}.$
@@
$\int e^{p_{0}|w-\bar{w}|}\leq C,$
where $\bar{w}=\frac{1}{|B_{1}|}\int w$. Take $p=p_{0}$ in (2.12). ∎
@@ 3.1. Introduction to fully nonlinear equations
Observe that the $L^{\epsilon}$-norm of a superharmonic function is controlled by its value at a point; hence, if we fix the value of $v$ at some point, say $v(0)=1$ without loss of generality, then $v$ belongs to $L^{\epsilon}$. See Lemma 4 and Lemma 7 below.
Conversely, the value of a subharmonic function is controlled by the nearby $L^{\epsilon}$-norm, so a fixed $L^{\epsilon}$-norm may govern the value at the center. See Lemma 1, Lemma 3 and Lemma 8 below.
@@ 3.3. Krylov-Safonov weak Harnack inequality
Let $v$ be a supersolution to (3.4) with $v(0)=1$. Then $v\in L^{\epsilon}_{weak}(B_{1/2})$, i.e.,
$|\{v>\lambda\}\cap B_{1/2}|\leq\lambda^{-\epsilon},\quad\forall\lambda>0.$
@@
First, we derive a contact estimate via the ABP method. To present a baby version, We claim that
@@
If $x_{0}\in\Gamma^{+}$, then $\tilde{v}(x_{0})<0$ (for more preciser auxiliary functions, seeing the sublemma below), which implies $v(x_{0})<2$. Hence $\Gamma^{+}\subset\{v<2\}$, and therefore
@@
sublemma 4 (Contact estimate).
@@
with constants $A,B$ chosen so that
@@
and set $\tilde{v}:=v+\gamma$ (see Figure 9). We can arrange $\gamma$ so that $\tilde{v}(0)=-2$ and $\inf\gamma\geq-M$, without loss of generality, set $M>2$.
@@
If $x\in\Gamma^{+}$, then $\tilde{v}(x)<0$ (it is clear in $B_{1/8}(x_{0})$; and $\Gamma^{+}$ can not escape from $B_{1/8}(x_{0})$ because $\tilde{v}$ is a supersolution outside $B_{1/8}(x_{0})$), which implies $v(x)<M$. Hence $\Gamma^{+}\subset\{v<M\}\cap B_{1/8}(x_{0})$, and the desired inequality follows. ∎
@@
which also solves (3.4). We consider two cases; if the first does not hold, we replace $w$ by 1w1-w. Recall that $w$ is a solution.
@@ 4.1. Concave assumption of $F$ and Gap
It remains to prove that $u\in C^{2,\alpha}$, i.e., $D^{2}u\in C^{\alpha}$. The $C^{\alpha}$ estimate (Corollary 4) requires a ”solution”, but we only know that $v$ is a subsolution. This gap constitutes the main difficulty in the Evans-Krylov theory. Fortunately, however, the relation
@@ 4.2. Deletion argument
Cover $\mathcal{H}(B_{1})$ by balls of radius $\delta$, which will be determined at the end of proof; their number satisfies $\lesssim\delta^{-n^{2}}$. Hence some ball $B_{\delta}(M_{3})$ satisfies the measure estimate
@@
For any $z\in\mathcal{H}^{-1}(B_{\delta}(M_{3}))$, we have $|u_{gg}(z)-u_{gg}(x_{3})|\lesssim\delta$. Choosing $\delta$ sufficiently small (universally) gives
$\mathbf{t}-u_{gg}(z)\gtrsim\frac{1}{8}-\delta\gtrsim\frac{1}{8},$
@@ 4.3.1. Krylov-Safonov
A simpler proof of the Krylov–Safonov theorem is provided in [mooney2019proof], whose approach does not rely on the localization technique.
4.3.2. Evans–Krylov
Since the Evans–Krylov theory requires concavity of $F$, a natural question is whether solutions are always $C^{2}$. The answer is affirmative for $n=2$ [nirenberg1953nonlinear] (see also Theorem 4.8 of [fernandez2023regularity]), where solutions are in fact $C^{2,\alpha}$. Counterexamples for $n\geq 5$ were constructed by Nadirashvili-Vladuts [nadirashvili2007nonclassical, nadirashvili2008singular, nadirashvili2013singular]. The cases $n=3,4$ remain open. We refer to [caffarelli1995fully] for a comprehensive treatment of fully nonlinear equations. Recently, Caffarelli-Yuan [caffarelli2000priori] established $C^{2,\alpha}$ regularity under a convexity assumption on the level set $\{F(D^{2}u)=0\}$; Cabré-Caffarelli [cabre2003interior] obtained the same results for certain nonconvex operators $F$. Later, Collins [collins2016c] and Goffi [goffi2024high] studied regularity under weaker convexity conditions.
4.3.3. Non-uniformly elliptic equations
The equations discussed in this note are all uniformly elliptic. In fact, many classical equations are locally uniformly elliptic but not globally uniformly elliptic, such as the Monge–Ampère equation $\mathrm{det}D^{2}u=0$ (or, more generally, the $\sigma_{k}$-Hessian equations), and the minimal surface equation $\mathrm{div}\left(\frac{\nabla u}{\sqrt{1+|\nabla u|^{2}}}\right)=0$ (or, more generally, the $\sigma_{k}$-curvature equations). The main difficulty in studying locally uniformly elliptic equations is that scaling can alter the uniform ellipticity. Consequently, the renormalization and blow-up arguments commonly used in previous proofs are no longer applicable.
Fortunately, to prove the weak Harnack inequality—for example, see Lemma 7—the so-called sliding paraboloids method, which estimates the contact measure by approximating a series of paraboloids, avoids the self-rescaling that would disturb uniform ellipticity. This method was first introduced by Cabré [cabre1997nondivergent], who extended the Krylov–Safonov estimate to manifolds with negative sectional curvature. Later, Savin [savin2009regularity], inspired also by Caffarelli-Córdoba [caffarelli1993elementary], used this approach to improve De Giorgi’s original flatness theorem [giusti1984minimal] (and, more generally, Allard regularity [allard1972first]), thereby proving the De Giorgi conjecture under a technical condition in dimensions $n\leq 8$. Furthermore, Savin [savin2007small] extended the idea to locally uniformly elliptic equations with small perturbations (or almost flatness), building on work of Caffarelli [caffarelli1987harnack] and Caffarelli–Wang [caffarelli1993harnack] (see further details in [caffarelli1995fully]). Subsequently, dos Prazeres and Teixeira [dos2016asymptotics] treated the non-homogeneous case under uniform ellipticity, while Fan [fan2025generalization] generalized it to the non-homogeneous case with only local uniform ellipticity. Wang [wang2013small] studied the parabolic case, and Yu [yu2017small] investigated the nonlocal case.